# Hierarchy-Boosted Funnel Learning for Identifying Semiconductors with Ultralow Lattice Thermal Conductivity

### v1

@@ Abstract
Data-driven machine learning (ML) has demonstrated tremendous potential in material property predictions. However, the scarcity of materials data with costly property labels in the vast chemical space presents a significant challenge for ML in efficiently predicting properties and uncovering structure-property relationships. Here, we propose a novel hierarchy-boosted funnel learning (HiBoFL) framework, which is successfully applied to identify semiconductors with ultralow lattice thermal conductivity ($\kappa_{\mathrm{L}}$). By training on only a few hundred materials targeted by unsupervised learning from a pool of hundreds of thousands, we achieve efficient and interpretable supervised predictions of ultralow $\kappa_{\mathrm{L}}$, thereby circumventing large-scale brute-force calculations without clear objectives. As a result, we provide a list of candidates with ultralow $\kappa_{\mathrm{L}}$ for potential thermoelectric applications and discover a new factor that significantly influences structural anharmonicity. This study offers a novel practical pathway for accelerating the discovery of functional materials.
@@ I Introduction
Emerging as a powerful technology of the data-driven paradigm in materials science, machine learning (ML) has considerably accelerated the design and discovery of promising materials in recent years,[1, 2, 3, 4] including ML interatomic potentials,[5] inverse design of materials,[6] efficient property predictions.[7, 8] Simultaneously, the advancements in high-performance computing greatly facilitate the establishment of diverse material-related databases utilizing density functional theory (DFT)-based high-throughput calculations (HTC), such as the Materials Project (MP),[9] Open Quantum Materials Database (OQMD),[10, 11] Automatic-FLOW for Materials Discovery (AFLOW),[12] Joint Automated Repository for Various Integrated Simulations (JARVIS),[13] etc. These continuously expanding databases also lay a solid foundation for the application of cutting-edge ML technology in the field of materials science.
@@
Figure 1: (a) Schematic of the novel HiBoFL framework, including data preparation, unsupervised learning, data annotation and supervised learning. (b) Workflow of applying HiBoFL framework to efficiently identify semiconductors with ultralow $\kappa_{\mathrm{L}}$.
@@ II.1 HiBoFL Framework for Accelerating the Discovery of Functional Materials
We then apply this HiBoFL framework to efficiently identify semiconductors with ultralow $\kappa_{\mathrm{L}}$ as shown in Figure 1b. In the first step, we obtain the material dataset from the MP database based on its application programming interface (API) and then apply a series of screening criteria to derive the first-level dataset for subsequent research. In the second step, chemical composition descriptors based on Magpie[39] and crystal structure descriptors derived from Voronoi tessellations[40] are used to featurize the materials in the first-level dataset. We use principal component analysis (PCA),[41] for dimensionality reduction thereby obtaining pivotal components. K-means clustering is then employed to identify materials with similar $\kappa_{\mathrm{L}}$, which is visualized in a low-dimensional space using t-distributed Stochastic Neighbor Embedding (t-SNE),[42] and problem-specific clusters are selected based on similarity design rules to form the second-level dataset. In the third step, we use the phonon-elasticity-thermal (PET) model[43] to perform low-cost HTC on the materials in the second-level dataset, establishing a local database based on the HTC results. On the one hand, a list of candidates for potential TE applications can be recommended and their statistical analysis can help us to summarize several domain knowledge of ultralow $\kappa_{\mathrm{L}}$. On the other hand, we can directly screen out candidate materials with ultralow $\kappa_{\mathrm{L}}$ for in-depth mechanism analysis, where the results are verified by accurately solving the Boltzmann transport equation and the phonon thermal transport mechanisms are further revealed. In the fourth step, we perform ensemble learning on the labeled local database for directly classifying ultralow $\kappa_{\mathrm{L}}$, training a robust CatBoost classifier[44]. The interpretable analysis of the pre-trained model based on the SHapley Additive exPlanations (SHAP) method[45, 46] further reveals the influence of important descriptors on ultralow $\kappa_{\mathrm{L}}$. A new factor capable of significantly reducing $\kappa_{\mathrm{L}}$ via enhancing structural anharmonicity is discovered, eventually building a bridge between the ML model interpretability and first-principles analysis.
@@ II.2 Preliminary High-Throughput Screening
The process of preliminary high-throughput screening to select materials for subsequent ML investigations is illustrated in Figure 2a. Initially, we start by acquiring all the materials from the MP database based on its API, resulting in a total of 154718 entries saved in a JSON file as the Python dictionary object. Of these, several specific criteria are set to narrow down the range of exploration. The first screening criterion focuses on the assessment of thermodynamic stability, in which the materials with energy above the convex hull ($E_{\mathrm{hull}}$, the formation energy difference between the target compound and its competing phases) no more than 0.05 eV/atom are considered to show a high likelihood of being synthesized in experiment. Further, the band gap ($E_{\mathrm{g}}$) can be directly retrieved from the MP database, set within the range of 0.1–2 eV for assessing electric conductivity. This specific interval is highly characteristic of semiconductors, demonstrating the inherent capacity of these screened materials for favorable electrical conductivity. Additionally, taking into account of the computational cost associated with $\kappa_{\mathrm{L}}$, our constraints on the material system involve ensuring that the number of atoms ($N_{\mathrm{atoms}}$) is less than 20 in the unit cell of a crystal structure and the number of elements ($N_{\mathrm{elements}}$) is below four in one compound, respectively. Ultimately, through excluding the materials containing hydrogen, lanthanides, and actinides, and conducting structural analysis, we further obtain 2675 three-dimensional (3D) crystal structures without calculation errors. These materials constitute the first-level dataset for our in-depth investigations.
@@ II.3 Unsupervised Learning for Identifying Materials with Similar $\kappa_{\mathrm{L}}$
We next carry out unsupervised learning to identify materials with a high likelihood of exhibiting relatively low $\kappa_{\mathrm{L}}$ based on similarity design rules. Following the generation of the first-level dataset as input, the materials should be transformed into the length-fixed vector as the so-called descriptors, which are the key point to distinguish different materials. Herein, we use two distinct types of descriptors to featurize these materials, namely, composition-based features and structure-based features, to account for the mapping relationships between different categories of descriptors and $\kappa_{\mathrm{L}}$ from the perspectives of chemical composition and crystal structure. Among them, composition-based features are generated based on Magpie data and denoted by {E}, including electronegativity, atomic number, fraction of electrons, etc., which are comprehensive enough to capture the characteristics of chemical compositions with different constituent elements and proportions. Through partitioning the crystal structures into Wigner-Seitz cells of each atom, a series of structure-based features can be derived from Voronoi tessellations based on the characteristics of local environment of each atom in the unit cell, which are denoted by {S}. In this manner, each material can generate a total of 273 descriptors automatically without any time-consuming DFT calculations based on its corresponding composition and structure, jointly denoted by {E, S}. Notably, these descriptors have been successfully applied to the ML predictions of $\kappa_{\mathrm{L}}$ at various temperatures, indicating their significant potential for mapping the property of $\kappa_{\mathrm{L}}$. After feature generation, it is crucial to preprocess all the data thereby enhancing the performance of clustering model. To address the large variations among different feature values and transform them to follow a normal distribution, both standardization and Quantile Transformer are employed for preprocessing these descriptors. Further, we perform PCA based on Singular Value Decomposition to project the preprocessed features to a lower dimensional space. PCA linearly combines original features into principal components (PCs), ensuring the hierarchical order based on their contributions, in which the first principal component captures the largest explainable variance, followed by the second one and so forth. From the curve of total explainable variance changing with the number of PCs as shown in Figure S1, we can conclude that 83 PCs suffice to account for 99% of the variance among all 2675 materials in the first-level dataset. Hence, these 83 PCs are extracted as the pivotal components for input into the following clustering algorithm.
@@
Subsequently, the k-means algorithm is utilized to identify the underlying associations with $\kappa_{\mathrm{L}}$ from these unlabeled materials, in which only one critical parameter is required to be predefined, i.e., the number of clusters k. Based on the analysis of the elbow method[47] and the silhouette coefficient[48] as depicted in Figure 2b, the elbow-like point of inflection on the curve emerges at a value of k = 7, indicating the location where inertia or distortion start decreasing significantly with a very slow rate, represents the optimal k value for the k-means clustering. At this stage, a relatively high silhouette coefficient also illustrate that each material exhibits strong similarity within its respective cluster, while different clusters are well-separated as much as possible. As a result, we partition the first-level dataset into seven clusters (from C1, C2, …, to C7), on the basis of similarities of these feature vectors derived from chemical compositions and crystal structures, in which the materials in the same cluster are considered to show a high likelihood of possessing similar $\kappa_{\mathrm{L}}$. Since the first two PCs only capture about 35% of the variance within entire data, it is insufficient for the clustering results to be intuitively visualized in a two-dimensional (2D) mapping. Unlike PCA, which emphasizes preserving large pairwise distances to maximize variance, t-SNE is a powerful non-linear dimensionality reduction technology for the visualization of high-dimensional data, aiming to maintain pairwise similarities among data points in a lower-dimensional space. We strictly use t-SNE only for the visualization of the resulting seven clusters of k-means algorithm as shown in Figure 3, projecting the 83 PCs into a 2D latent space comprised of two t-SNE components. The result intuitively demonstrates the high-quality clustering with distinct separation between each cluster, where each point represents a material and its color relates to the corresponding category of cluster.
@@
Figure 4: (a–d) Statistical analysis of HTC results in the second-level dataset. (a) Pie chart of $\kappa_{\mathrm{PET}}$ separated by a threshold of 2 W/mK and distribution of different formula types with counts exceeding 20. (b–d) Distribution of the materials with $\kappa_{\mathrm{PET}}$ no greater than 2 W/mK: (b) Heat map over the periodic table of elements for easy visualization of element counts. (c) Box plot of different crystal systems. (d) Density scatter plot of shear modulus and bulk modulus. (e) Calculated $\kappa_{\mathrm{L}}$ as the function of temperature at different axes in Cs2SnSe3 and Cs2GeSe3 by solving the phonon Boltzmann transport equation.
@@ II.4 HTC of Specific Clusters and Statistical Analysis
Given the success of unsupervised learning in significantly reducing the broad material search space, it has become feasible to further extend the second-level dataset into a labeled repository of thermal conductivity through HTC at affordable computational costs. Based on the PET empirical equation within the high-throughput framework, we derive the $\kappa_{\mathrm{PET}}$ values at 300 K ignoring the anisotropy for materials in the second-level dataset, while excluding structures that do not satisfy the mechanical stability criteria. The basic information and corresponding $\kappa_{\mathrm{PET}}$ values of the resulting 661 materials are then stored in a local database using MongoDB,[60] serving as a valuable repository for researchers to retrieve data and prioritize detailed theoretical and experimental investigations. Of particular note is that nearly 70% of these semiconductors exhibit the $\kappa_{\mathrm{PET}}$ values no greater than 2 W/mK (Figure 4a), with a considerable portion of these materials remaining unreported to date. These materials also constitute a list of candidate materials with potential applications in the TE field, which undoubtedly prove that the specific clusters we identified through unsupervised learning indeed contain a significant number of materials with low thermal conductivity.
@@
In addition, we also count the distribution of different formula types with counts exceeding 20 in these materials. Apparently, the structures represented by the two types of formula anonymous dominate in quantity, namely the ABC2 type characterized by the diamond-like structure, and the ABC3 type represented by the perovskite-type structure. To gain a more intuitive insight into the distribution patterns of materials with low thermal conductivity, we plot a heat map over the periodic table of elements in Figure 4b, illustrating the count of elements present in these compounds with $\kappa_{\mathrm{PET}}$ no greater than 2 W/mK. Sulfur, selenium, tellurium, and oxygen belonging to the chalcogens consecutively occupy the largest counts among the anion elements, in which sulfur is the most abundant with a count of 102. As for the cation elements, silver, cesium, potassium, copper, and rubidium occupy the highest abundance respectively. This can be explained by the fact that materials composed of heavy elements or characterized by weak chemical bonding typically exhibit lower $\kappa_{\mathrm{L}}$. Notably, a few previous studies have constrained the search space of materials within the above-mentioned characteristics to investigate those with low thermal conductivity, such as high-throughput screening in chalcogenide ABC3 perovskites[61] or diamond-like ABC2 compounds,[62] detailed analysis in the IV–VI chalcogenides[63] and so forth.[64, 65, 66, 38] The box plot of different crystal systems indicates that the $\kappa_{\mathrm{PET}}$ values of materials in each crystal system primarily cluster around 0.5 W/mK, with the monoclinic system exhibiting the largest quantity (Figure 4c). We also show the density distributions of the shear modulus and bulk modulus obtained by the Voigt–Reuss–Hill (VRH) method as shown in Figure 4d, which are primarily concentrated within the relatively small range of $\sim$50 GPa. This suggests that materials with lower shear and bulk moduli might be more likely to exhibit lower thermal conductivity. These statistical analyses provide deep insights into the regulation of $\kappa_{\mathrm{L}}$, which can offer prior domain knowledge for researchers in selecting specific systems to obtain ultralow $\kappa_{\mathrm{L}}$.
@@
To further validate the results and conduct in-depth analysis of the phonon thermal transport mechanism, we calculate more precise $\kappa_{\mathrm{L}}$ values according to first-principles derived force constants and Boltzmann transport theory for the unreported materials Cs2SnSe3 and Cs2GeSe3, which are screened out based on the formula type ranking among the materials with the lowest $\kappa_{\mathrm{PET}}$ values. Figure 4e depicts the calculated $\kappa_{\mathrm{L}}$ as the function of temperature ranging from 100 to 500 K at different axes in the discussed semiconductors, in which all the intrinsic $\kappa_{\mathrm{L}}$ values show obvious anisotropy and gradually decrease following the T–1 manner with the temperature increasing just as the hallmark of Umklapp scattering. The rise in temperature results in an elevation in the equilibrium phonon population, consequently leading to intense phonon-phonon collisions, as delineated by $\overline{n}\approx{k_{\mathrm{B}}T}/{\hbar\omega}\ (T\gg\Theta_{\mathrm{D}})$, where $\overline{n}$ is the average number of phonons, $k_{\mathrm{B}}$ is the Boltzmann constant, T is the temperature, $\omega$ is the frequency, $\hbar$ is the reduced Planck constant and $\Theta_{\mathrm{D}}$ is the Debye temperature. Moreover, these materials all exhibit intrinsically ultralow $\kappa_{\mathrm{L}}$, with values all below 0.25 W/mK in any direction at 300 K, significantly lower than the $\kappa_{\mathrm{L}}$ ($\sim$2.3 W/mK) of the traditional TE material PbTe.[67] The results indicate the potential application of Cs2SnSe3 and Cs2GeSe3 in the TE field, further substantiating the effectiveness of our previously adopted approach—combining unsupervised learning with HTC to discover semiconductors with ultralow $\kappa_{\mathrm{L}}$.

### v2

@@ Abstract
Data-driven machine learning (ML) has demonstrated tremendous potential in material property predictions. However, the scarcity of materials data with costly property labels in the vast chemical space presents a significant challenge for ML in efficiently predicting properties and uncovering structure-property relationships. Here, we propose a novel hierarchy-boosted funnel learning (HiBoFL) framework, which is successfully applied to identify semiconductors with ultralow lattice thermal conductivity ($\kappa_{\mathrm{L}}$). By training on only a few hundred materials targeted by unsupervised learning from a pool of hundreds of thousands, we achieve efficient and interpretable supervised predictions of ultralow $\kappa_{\mathrm{L}}$, thereby circumventing large-scale brute-force ab initio calculations without clear objectives. As a result, we provide a list of candidates with ultralow $\kappa_{\mathrm{L}}$ for potential thermoelectric applications and discover a new factor that significantly influences structural anharmonicity. This HiBoFL framework offers a novel practical pathway for accelerating the discovery of functional materials.
@@ I Introduction
Emerging as a powerful technology of the data-driven paradigm in materials science, machine learning (ML) has considerably accelerated the design and discovery of promising materials in recent years,[1, 2, 3, 4, 5] including ML interatomic potentials,[6] inverse design of materials,[7] efficient property predictions.[8, 9] Simultaneously, the advancements in high-performance computing greatly facilitate the establishment of diverse material-related databases utilizing density functional theory (DFT)-based high-throughput calculations (HTC), such as the Materials Project (MP),[10] Open Quantum Materials Database (OQMD),[11, 12] Automatic-FLOW for Materials Discovery (AFLOW),[13] Joint Automated Repository for Various Integrated Simulations (JARVIS),[14] etc. These continuously expanding databases also lay a solid foundation for the application of cutting-edge ML technology in the field of materials science.
@@
Figure 1: (a) Schematic of the HiBoFL framework, including data preparation, unsupervised learning, data annotation and supervised learning. (b) Workflow of applying the HiBoFL framework to efficiently identify semiconductors with ultralow $\kappa_{\mathrm{L}}$.
@@ II.1 HiBoFL Framework for Accelerating the Discovery of Functional Materials
We then apply this HiBoFL framework to efficiently identify semiconductors with ultralow $\kappa_{\mathrm{L}}$ as shown in Figure 1b. In the first step, we obtain the material dataset from the MP database based on its application programming interface (API) and then apply a series of screening criteria to derive the first-level dataset for subsequent research. In the second step, chemical composition descriptors based on Magpie[40] and crystal structure descriptors derived from Voronoi tessellations[41] are used to featurize the materials in the first-level dataset. We use principal component analysis (PCA),[42] for dimensionality reduction thereby obtaining pivotal components. K-means clustering[43, 44] is then employed to identify materials with similar $\kappa_{\mathrm{L}}$, which is visualized in a low-dimensional space using t-distributed stochastic neighbor embedding (t-SNE),[45] and problem-specific clusters are selected based on similarity design rules to form the second-level dataset. In the third step, we use the phonon-elasticity-thermal (PET) model[46] to perform low-cost HTC on the materials in the second-level dataset, establishing a local database based on the HTC results. On the one hand, a list of candidates for potential TE applications can be recommended and their statistical analysis can help us to summarize several domain knowledge of ultralow $\kappa_{\mathrm{L}}$. On the other hand, we can directly screen out candidate materials with ultralow $\kappa_{\mathrm{L}}$ for in-depth mechanism analysis, where the results are verified by accurately solving the phonon Boltzmann transport equation (BTE) and the phonon thermal transport mechanisms are further revealed. In the fourth step, we perform ensemble learning on the labeled local database for directly classifying ultralow $\kappa_{\mathrm{L}}$, training a robust CatBoost classifier[47]. The interpretable analysis of the pre-trained model based on the SHapley Additive exPlanations (SHAP) method[48, 49, 50] further reveals the influence of important descriptors on ultralow $\kappa_{\mathrm{L}}$. A new factor capable of significantly reducing $\kappa_{\mathrm{L}}$ via enhancing structural anharmonicity is discovered, eventually building a bridge between the ML model interpretability and first-principles analysis.
@@ II.2 Preliminary High-Throughput Screening
The process of preliminary high-throughput screening to select materials for subsequent ML investigations is illustrated in Figure 2a. Initially, we start by acquiring all the materials from the MP database based on its API, resulting in a total of 154718 entries saved in a JSON file as the Python dictionary object. Of these, several specific criteria are set to narrow down the range of exploration. The first screening criterion focuses on the assessment of thermodynamic stability, in which the materials with energy above the convex hull ($E_{\mathrm{hull}}$, the formation energy difference between the target compound and its competing phases) no more than 0.05 eV/atom are considered to show a high likelihood of being synthesized in experiments. Further, the band gap ($E_{\mathrm{g}}$) can be directly retrieved from the MP database, set within the range of 0.1–2 eV for assessing electric conductivity. This specific interval is highly characteristic of semiconductors, demonstrating the inherent capacity of these screened materials for favorable electrical conductivity. Additionally, taking into account of the computational cost associated with $\kappa_{\mathrm{L}}$, our constraints on the material system involve ensuring that the number of atoms ($N_{\mathrm{atoms}}$) is less than 20 in the unit cell of a crystal structure and the number of elements ($N_{\mathrm{elements}}$) is below four in one compound, respectively. Ultimately, through excluding the materials containing hydrogen, lanthanides, and actinides, and conducting structural analysis, we further obtain 2675 three-dimensional (3D) crystal structures without calculation errors. These materials constitute the first-level dataset for our in-depth investigations.
@@ II.3 Unsupervised Learning for Identifying Materials with Similar $\kappa_{\mathrm{L}}$
We next carry out unsupervised learning to identify materials with a high likelihood of exhibiting relatively low $\kappa_{\mathrm{L}}$ based on similarity design rules. Following the generation of the first-level dataset as input, the materials should be transformed into the length-fixed vector as the so-called descriptors, which are the key point to distinguish different materials. Herein, we use two distinct types of descriptors to featurize these materials, namely, composition-based features and structure-based features, to account for the mapping relationships between different categories of descriptors and $\kappa_{\mathrm{L}}$ from the perspectives of chemical composition and crystal structure. Among them, composition-based features are generated based on Magpie data and denoted by {E}, including electronegativity, atomic number, fraction of electrons, etc., which are comprehensive enough to capture the characteristics of chemical compositions with different constituent elements and proportions. Through partitioning the crystal structures into Wigner-Seitz cells of each atom, a series of structure-based features can be derived from Voronoi tessellations based on the characteristics of the local environment of each atom in the unit cell, which are denoted by {S}. In this manner, each material can generate a total of 273 descriptors automatically without any time-consuming DFT calculations based on its corresponding composition and structure, jointly denoted by {E, S}. Notably, these descriptors have been successfully applied to the ML predictions of $\kappa_{\mathrm{L}}$ at various temperatures, indicating their significant potential for mapping the property of $\kappa_{\mathrm{L}}$. After feature generation, it is crucial to preprocess all the data thereby enhancing the performance of clustering model. To address the large variations among different feature values and transform them to follow a normal distribution, both standardization and quantile transformer are employed for preprocessing these descriptors. Further, we perform PCA based on singular value decomposition to project the preprocessed features to a lower dimensional space. PCA linearly combines original features into principal components (PCs), ensuring the hierarchical order based on their contributions, in which the first principal component captures the largest explainable variance, followed by the second one and so forth. From the curve of total explainable variance changing with the number of PCs as shown in Figure S1, we can conclude that 83 PCs suffice to account for 99% of the variance among all 2675 materials in the first-level dataset. Hence, these 83 PCs are extracted as the pivotal components for input into the following clustering algorithm.
@@
Subsequently, the k-means algorithm is utilized to identify the underlying associations with $\kappa_{\mathrm{L}}$ from these unlabeled materials, in which only one critical parameter is required to be predefined, i.e., the number of clusters k.[51] Based on the analysis of the elbow method[52] and the silhouette coefficient[53] as depicted in Figure 2b, the elbow-like point of inflection on the curve emerges at a value of k = 7, indicating the location where inertia or distortion starts decreasing significantly at a very slow rate, represents the optimal k value for the k-means clustering. At this stage, a relatively high silhouette coefficient also illustrates that each material exhibits strong similarity within its respective cluster, while different clusters are well-separated as much as possible. As a result, we partition the first-level dataset into seven clusters (from C1, C2, …, to C7), on the basis of similarities of these feature vectors derived from chemical compositions and crystal structures, in which the materials in the same cluster are considered to show a high likelihood of possessing similar $\kappa_{\mathrm{L}}$. Since the first two PCs only capture about 35% of the variance within entire data, it is insufficient for the clustering results to be intuitively visualized in a two-dimensional (2D) mapping. Unlike PCA, which emphasizes preserving large pairwise distances to maximize variance, t-SNE is a powerful non-linear dimensionality reduction technology for the visualization of high-dimensional data, aiming to maintain pairwise similarities among data points in a lower-dimensional space. We strictly use t-SNE only for the visualization of the resulting seven clusters of k-means algorithm as shown in Figure 3, projecting the 83 PCs into a 2D latent space comprised of two t-SNE components. The result intuitively demonstrates the high-quality clustering with distinct separation between each cluster, where each point represents a material and its color relates to the corresponding category of cluster.
@@
Figure 4: (a–d) Statistical analysis of HTC results in the second-level dataset. (a) Pie chart of $\kappa_{\mathrm{PET}}$ separated by a threshold of 2 W/mK and distribution of different formula types with counts exceeding 20. (b–d) Distribution of the materials with $\kappa_{\mathrm{PET}}$ no greater than 2 W/mK: (b) Heat map over the periodic table of elements for easy visualization of element counts. (c) Box plot of different crystal systems. (d) Density scatter plot of shear modulus and bulk modulus. (e) Calculated $\kappa_{\mathrm{L}}$ as the function of temperature at different axes in Cs2SnSe3 and Cs2GeSe3 by solving the phonon BTE.
@@ II.4 HTC of Specific Clusters and Statistical Analysis
Given the success of unsupervised learning in significantly reducing the broad material search space, it has become feasible to further extend the second-level dataset into a labeled repository of thermal conductivity through HTC at affordable computational costs. Based on the PET empirical equation within the high-throughput framework, we derive the $\kappa_{\mathrm{PET}}$ values at 300 K ignoring the anisotropy for materials in the second-level dataset, while excluding structures that do not satisfy the mechanical stability criteria. The basic information and corresponding $\kappa_{\mathrm{PET}}$ values of the resulting 661 materials are then stored in a local database using MongoDB,[65] serving as a valuable repository to retrieve data and prioritize detailed theoretical and experimental investigations. Of particular note is that nearly 70% of these semiconductors exhibit the $\kappa_{\mathrm{PET}}$ values no greater than 2 W/mK (Figure 4a), with a considerable portion of these materials remaining unreported to date. These materials also constitute a list of candidate materials with potential applications in the TE field (Table S2), which undoubtedly prove that the specific clusters we identified through unsupervised learning indeed contain a significant number of materials with low thermal conductivity.
@@
In addition, we also count the distribution of different formula types with counts exceeding 20 in these materials. Apparently, the structures represented by the two types of formula anonymous dominate in quantity, namely the ABC2 type characterized by the diamond-like structure, and the ABC3 type represented by the perovskite-type structure. To gain a more intuitive insight into the distribution patterns of materials with low thermal conductivity, we plot a heat map over the periodic table of elements in Figure 4b, illustrating the count of elements present in these compounds with $\kappa_{\mathrm{PET}}$ no greater than 2 W/mK. Sulfur, selenium, tellurium, and oxygen belonging to the chalcogens consecutively occupy the largest counts among the anion elements, in which sulfur is the most abundant with a count of 102. As for the cation elements, silver, cesium, potassium, copper, and rubidium occupy the highest abundance, respectively. This can be explained by the fact that materials composed of heavy elements or characterized by weak chemical bonding typically exhibit lower $\kappa_{\mathrm{L}}$. Notably, a few previous studies have constrained the search space of materials within the above-mentioned characteristics to investigate those with low thermal conductivity, such as high-throughput screening in chalcogenide ABC3 perovskites[66] or diamond-like ABC2 compounds,[67] detailed analysis in the IV–VI chalcogenides[68] and so forth.[69, 70, 71, 39] The box plot of different crystal systems indicates that the $\kappa_{\mathrm{PET}}$ values of materials in each crystal system primarily cluster around 0.5 W/mK, with the monoclinic system exhibiting the largest quantity (Figure 4c). We also show the density distributions of the shear modulus and bulk modulus obtained by the Voigt–Reuss–Hill (VRH) method as shown in Figure 4d, which are primarily concentrated within the relatively small range of $\sim$50 GPa. This suggests that materials with lower shear and bulk moduli might be more likely to exhibit lower thermal conductivity. These statistical analyses provide deep insights into the regulation of $\kappa_{\mathrm{L}}$, which can offer prior domain knowledge for researchers in selecting specific systems to obtain ultralow $\kappa_{\mathrm{L}}$.
@@
To further validate the results and conduct in-depth analysis of the phonon thermal transport mechanism, we calculate more precise $\kappa_{\mathrm{L}}$ values according to first-principles derived force constants and Boltzmann transport theory for the unreported materials Cs2SnSe3 and Cs2GeSe3, which are screened out based on the formula type ranking among the materials with the lowest $\kappa_{\mathrm{PET}}$ values. Figure 4e depicts the calculated $\kappa_{\mathrm{L}}$ as the function of temperature ranging from 100 to 500 K at different axes in the discussed semiconductors, in which all the intrinsic $\kappa_{\mathrm{L}}$ values show obvious anisotropy and gradually decrease following the T–1 manner with the temperature increasing just as the hallmark of Umklapp scattering. The rise in temperature results in an elevation in the equilibrium phonon population, consequently leading to intense phonon-phonon collisions, as delineated by $\overline{n}\approx{k_{\mathrm{B}}T}/{\hbar\omega}\ (T\gg\Theta_{\mathrm{D}})$, where $\overline{n}$ is the average number of phonons, $k_{\mathrm{B}}$ is the Boltzmann constant, T is the temperature, $\omega$ is the phonon frequency, $\hbar$ is the reduced Planck constant and $\Theta_{\mathrm{D}}$ is the Debye temperature. Moreover, these materials all exhibit intrinsically ultralow $\kappa_{\mathrm{L}}$, with values all below 0.25 W/mK in any direction at 300 K, significantly lower than the $\kappa_{\mathrm{L}}$ ($\sim$2.3 W/mK) of the traditional TE material PbTe.[72] The results indicate the potential application of Cs2SnSe3 and Cs2GeSe3 in the TE field, further substantiating the effectiveness of our previously adopted approach—combining unsupervised learning with HTC to discover semiconductors with ultralow $\kappa_{\mathrm{L}}$.